Each of the following mixtures is placed in a dialyzing bag and immersed in distilled water. Which substances will be found outside the bag in the distilled water?

starch solution (colloid) and alanine, an amino acid, solution

starch solution (colloid) and alanine, an amino acid, solution

starch
alanine
starch and alanine
none of the above

NaCl solution and starch solution (colloid)

solution and starch solution (colloid)

starch
Na+ and Cl− ions
starch and Na+ and Cl− ions

none of the above

urea solution

urea solution

urea hydrochloride
urea hydroxide
urea

none of the above

A lactated Ringer's solution with 5% glucose contains Na+ 134 mEq/L , K+  4 mEq/L , Ca2+  2 mEq/L , Cl−  110. mEq/L , lactate−  30. mEq/L , and 50. g/L glucose (C6H12O6).

What is the total positive charge of the electrolytes?

Express your answer in milliequivalents per liter as an integer.

What is the total negative charge of the electrolytes?

Express your answer in milliequivalents per liter as an integer.

What is the osmolarity (Osm) of the solution?

Express the osmolarity to two decimal places.
Answer #1

The total positive charge of the electrolytes = 134 + 4 + 2 = 140 mEq/L

The total negative charge of the electrolytes = 110 + 30 = 140 mEq/L

The osmolarity of the solution = 140.00 mEq/L